A110 inch pipe is cut into two pieces one piece is four times the length of the other find the length of the two pieces?
The short piece is ..... inches long.
The long piece is ..... inches long.

Answer by waynest(281)

1st piece = x
2nd piece = 4x
x + 4x = 110
5x = 110
5x/5 = 110/5
x = 22
1st piece = 22 inches
2nd piece = 22 * 4
= 88